# On an approach to canonicalizing elliptic Feynman integrals

### v1

@@ Abstract
We present a systematic method for the construction of canonical bases for univariate elliptic Feynman integrals with multiple kinematic scales, which frequently arise in phenomenologically relevant scattering processes. The construction is performed in the Legendre normal form of elliptic curves, where the geometric propoerties of the curves are simple and explicit and further kinemtic singularities are present as marked points. The canonical bases are constructed using Abelian differentials of three kinds with a universal linear transformation. The bases constructed in the normal form can be mapped to generic integral families via an appropriate Möbius transformation. As a demonstration, we discuss the application of our method to several concrete examples, where we show that the $\varepsilon$-factorization of sub-sector dependence can also be simply done. Our method can be readily applied to more complicated integral families straightforwardly.
@@ 1 Introduction
The differential equations method Kotikov:1990kg ; Kotikov:1991pm ; Remiddi:1997ny ; Gehrmann:1999as , and specifically the canonical differential equations approach Henn:2013pwa , represents the state-of-the-art method for calculating MIs analytically. This approach leverages an $\varepsilon$-factorized form, allowing MIs to be expressed as Laurent expansions in the dimensional regulator $\varepsilon$, with Chen’s iterated integrals Chen:1977oja as coefficients. [^1] A canonical basis has other nice properties in a addition to $\varepsilon$-factorization. The simplest iterated integrals found in Feynman integrals are multiple polylogarithms (MPLs) Goncharov:1998kja ; Goncharov:2001iea ; Vollinga:2004sn , which are iterated integrals over the genus-00 Riemann sphere with marked points, or say over moduli space ${\cal M}_{0,n}$ Bogner:2014mha . For MPL integral families, canonical differential equations can be achieved through rational and algebraic transformations alone, and several tools are available to assist in identifying such transformations Gituliar:2017vzm ; Prausa:2017ltv ; Meyer:2017joq ; Lee:2020zfb ; Dlapa:2020cwj . An alternative approach constructs canonical bases through ${\rm d}\log$-forms Henn:2020lye ; Chen:2020uyk ; Chen:2022lzr . [^2] However, not all integrals in ${\rm d}\log$-forms are polylogarithmic, see Duhr:2020gdd . MPLs exhibit properties that facilitate algorithmic extraction of critical information, aiding in the construction of canonical bases and enriching our understanding of their mathematical structure.
@@
Due to the success of the method of canonical differential equations in solving the MPL integral families, there have been substantial efforts to extend the method to the elliptic cases appearing in various scattering processes Muller:2022gec ; Jiang:2023jmk ; Gorges:2023zgv ; Delto:2023kqv ; Wang:2024ilc ; Forner:2024ojj ; Schwanemann:2024kbg ; Marzucca:2025eak ; Becchetti:2025rrz ; Becchetti:2025oyb . In such cases, there are subtleties in the precise definition of “canonical” Broedel:2018iwv ; Broedel:2018qkq ; Broedel:2019hyg ; Dlapa:2022wdu ; Frellesvig:2023iwr ; Duhr:2024uid ; Duhr:2025lbz , but one may at least attempt to obtain an $\varepsilon$-factorized system of differential equations which still offers simplification in the calculations. There are now several methods for deriving $\varepsilon$-factorized bases for elliptic cases vonManteuffel:2017hms ; Adams:2018bsn ; Frellesvig:2021hkr ; Pogel:2022vat ; Giroux:2022wav ; Dlapa:2022wdu ; Gorges:2023zgv , and some of them can even be extended to more complex geometries, such as Calabi-Yau varieties. At the moment, finding $\varepsilon$-factorized bases still requires non-trivial case-by-case analyses, especially when there are multiple kinematic marked points Bogner:2019lfa ; Gorges:2023zgv ; Delto:2023kqv ; Schwanemann:2024kbg ; Becchetti:2025rrz . This raises the natural question of whether it is possible to directly construct the $\varepsilon$-factorized basis for an elliptic Feynman integral family analogous to the approach for MPLs in Chen:2020uyk ; Chen:2022lzr .
@@
In this paper, we introduce a novel method to systematize the construction of canonical bases for univariate elliptic Feynman integrals under the maximal cut. We use the term “canonical” here since it meets most of the important properties advocated by the literature Dlapa:2022wdu ; Frellesvig:2023iwr ; Duhr:2024uid . Our method utilizes Legendre normal form of elliptic curves, where the geometric properties of the curves are simple and explicit. [^3] The Legendre normal form has also be employed in Broedel:2018rwm . Similarly, the Rosenhain normal form for the hyperelliptic case has been explored recently in Duhr:2024uid . We then construct the pre-canonical basis using Abelian differentials of three kinds, accounting for an arbitrary number of kinematic marked points. The differential equations of the pre-canonical basis can be straightforwardly converted to the $\varepsilon$-factorized form with a universal transformation. Such a construction can be applied to arbitrary elliptic curves with appropriate Möbius transformations which transform the curves to the Legendre normal form.
@@
To demonstrate our construction, we apply the method to several concrete examples: fully massive sunrise and non-planar double box diagrams, each with two kinds of mass configurations. We show that it is rather straightforward to derive an $\varepsilon$-factorized system of differential equations within the top sectors. To obtain a canonical basis for the entire integral family, we still need to address the non-$\varepsilon$-factorized dependence on the sub-sectors. Ideally speaking, one would imagine that it is possible to construct the full integrand without any cut [^4] We have demonstrated an example with the next-to-maximal cut in Sec. 3.2., combining the top-sector construction in this work and the ${\rm d}\log$-form construction in Chen:2020uyk ; Chen:2022lzr . In this work, we simply convert the differential equations to the $\varepsilon$-factorized form through linear transformation. We show in the examples that the nice properties of our basis helps a lot in this process.
@@
This paper is organized as follows. In Sec. 2, We present the construction of the canonical basis under maximal cut, which involves the construction in the Legendre normal form and the transformation to an elliptic curve defined by an arbitrary degree-4 polynomial. We demonstrate the procedure with several examples in Sec. 3, where the transformation to a canonical basis for the full family without cut is discussed. The summary and outlook are provided in Sec. 4. There are a few appendices at the end of the paper. A different strategy vonManteuffel:2017hms ; Frellesvig:2021hkr for deriving an $\varepsilon$-factorized basis is presented in App. A. The construction using the Jacobi normal form of elliptic curves is discussed in App. B. Some lengthy expressions relevant for the examples in Sec. 3 are collected in App. C. Special relations between elliptic integrals are compiled in App. D.
@@ 2.1 Basics for elliptic Feynman integrals
$I_{\nu_{1}\cdots\nu_{N}}=e^{L\varepsilon\gamma_{E}}\left(\mu^{2}\right)^{\nu-Ld/2}\int\,\bigg{[}\,\prod_{j=1}^{L}\frac{{\rm d}^{d}k_{j}}{i\pi^{d/2}}\,\bigg{]}\frac{1}{D_{1}^{\nu_{1}}\cdots D_{N}^{\nu_{N}}}\,,$ (1)
@@
$u=u(z)=\big{(}P_{4}\left(z\right)\big{)}^{-1/2}\prod_{i=1}^{n}\left(z-c_{i}\right)^{-\beta_{i}\varepsilon}\,,$ (3)
@@ 2.2 The construction of the pre-canonical basis
$u_{L}(x)=P_{L}(x)^{-1/2}\prod_{i=2}^{n+1}(x-e_{i})^{-\beta_{i}\varepsilon}\,,$ (13)
@@
${\rm d}\vec{I}=\bm{A}\,\vec{I}=\left({\bm{A}^{(0)}}+\varepsilon{\bm{A}^{(1)}}\right)\vec{I}\,,$ (19)
@@ 3.1.1 Definition
$u(z_{4})=P_{4}(z_{4})^{-1/2}\prod_{i=1}^{5}(z_{4}-c_{i})^{-\beta_{i}\varepsilon}\,.$ (62)
@@ 3.2.1 Definition
$s\equiv\left(p_{1}+p_{2}\right)^{2},\quad t\equiv\left(p_{1}+p_{3}\right)^{2}.$ (75)
@@
$u(z_{6},z)=P_{4}(z)^{-1/2}P_{2}(z_{6},z)^{-1/2}\prod_{i=1}^{4}(z-c_{i})^{-\beta_{i}\varepsilon}\prod_{j=6}^{7}(z_{6}-c_{j})^{-\beta_{j}\varepsilon}\,.$ (79)
@@ 3.2.3 Mixing with non-elliptic sectors
The $\varepsilon^{0}$-terms can be removed with regular “integrating out” strategy, and it turns out there are no new functions needed in the transformation beyond the periods $\varpi_{0}(\lambda)$ and $\vartheta_{0}(e_{5},\lambda)$ in ${\cal T}_{\rm can}$. For simplicity, we suppress the explicit expressions here and they can be found in ancillary files.
@@ 4 Conclusion and outlook
In this paper, we develop a systematic method to construct a canonical basis for elliptic Feynman integral families involving multiple kinematic scales, which exhibit a univariate integral in the Baikov representation under maximal cut. The core idea is to perform the construction using the Legendre normal form of elliptic curves, where the geometric information is explicit, and the apply a Möbius transformation to map it back to the original integral family. The pre-canonical basis integrals under maximal cut are expressed as linear combinations of Abelian differentials, without elliptic integrals in the combination coefficients. This feature provides simplification when performing the $\varepsilon$-factorization for the dependence on sub-sectors. Furthermore, the canonical bases exhibit favorable behaviors around chosen degenerate points, where they asymptotically take the ${\rm d}\log$-forms. This helps to obtain simple boundary conditions for the differential equations. Around the degenerate point, only the dependence of the basis integrals on the marked points becomes significant, allowing the integrals to be expressed in terms of MPLs in this context.
@@
There can be different ways in the construction of the pre-canonical bases and the transformations bringing them to the canonical ones. Besides the method described in the main text, we have explored several other possibilities which are described in the appendices. In App. A, we show that our pre-canonical basis can be transformed to a canonical one by converting the period matrix into a constant matrix without $\varepsilon$-dependent transformation vonManteuffel:2017hms ; Frellesvig:2021hkr . Despite a simpler $\varepsilon$-dependence, this method introduces numerous elliptic integrals into the transformation, which, in turn, appear in the $\varepsilon$-factorized differential equations. In addition, this approach inherently leads to the transformation with logarithmic behaviors near degenerate points due to the complementary arguments $\lambda$ and $1-\lambda$ in elliptic integrals of the first and second kinds. Relevant discussions are included in Frellesvig:2023iwr . In contrast, the approach that we take avoids these issues, and the resulting canonical bases satisfy most of the nice properties advocated in Dlapa:2022wdu ; Frellesvig:2023iwr ; Duhr:2024uid .
@@
To demonstrate our method, we use fully massive sunrise families and the non-planar double-box families with an internal mass as a proof of concept. Our method is similar to those in vonManteuffel:2017hms ; Adams:2018bsn ; Frellesvig:2021hkr ; Giroux:2022wav ; Dlapa:2022wdu ; Pogel:2022vat ; Gorges:2023zgv , but provides a more streamline framework for the construction. With the universal construction in the normal form, the results can be readily applied to generic elliptic integral families. Moreover, since the proposed approach is not inherently limited to elliptic cases, it is expected to extend to more complex geometries, e.g., hyperelliptic cases, as explored in the recent work Duhr:2024uid . For example, using the canonical basis integrals constructed for the four-parameter Lauricella functions, combined with an appropriate Möbius transformation, one may derive the canonical basis for the equal-mass non-planar double (crossed) box family, at least under the maximal cut. One can also try constructing a canonical basis in suitable (Rosenhain) normal form family with kinematic marked points, which can be potentially applied for generic hyperelliptic (of fixed genus) Feynman integrals, with suitable Möbius transformations.
@@
Our method can be applied to integral families where multiple sectors involve the same elliptic curve. In such cases, the construction can be done sector-by sector. However, we have found that it is helpful to perform the construction for these sectors together by loosening the maximal cut. This has been demonstrated for the non-planar double box family in Sec. 3.2, where both the top-sector and one of the triangle sectors are elliptic. It remains possible that one may construct the full integrand without any cut, combining the elliptic-sector construction in this work and the ${\rm d}\log$-form construction in Chen:2020uyk ; Chen:2022lzr . It is also interesting to explore the application of our method to cases where multiple elliptic curves are involved. In this case, there can be non-trivial interplays among the Abelian differentials in different elliptic sectors, and it is interesting to see how these can be handled within our approach. Another complication arises in the boundary conditions, since the degenerate points of the different elliptic curves may not coincide. It remains to be seen whether one may utilize the degenerate information separately, and connect them via the differential equations. We leave these for future investigations.

### v2

@@ Abstract
We develop a systematic approach to systematically constructing canonical bases for univariate elliptic Feynman integrals with multiple kinematic scales, which frequently arise in phenomenologically relevant scattering processes. The construction is performed in the Legendre normal form of elliptic curves, where the geometric properties of the curves are simple and explicit, and further kinematic singularities are present as marked points. The simplicity of the normal form facilitates the construction of universal canonical bases with an arbitrary number of marked points, which can be mapped to generic univariate elliptic integral families via an appropriate Möbius transformation. Therefore, our result can be readily applied to these cases straightforwardly. As a demonstration, we discuss the application of our method to several concrete examples, where we show that the $\varepsilon$-factorization of sub-sector dependence can also be simply done.
@@ 1 Introduction
The differential equations method Kotikov:1990kg ; Kotikov:1991pm ; Remiddi:1997ny ; Gehrmann:1999as , and specifically the canonical differential equations approach Henn:2013pwa , represents the state-of-the-art method for calculating MIs analytically. This approach leverages an $\varepsilon$-factorized form, allowing MIs to be expressed as Laurent expansions in the dimensional regulator $\varepsilon$, with Chen’s iterated integrals Chen:1977oja as coefficients. [^1] A canonical basis has other nice properties in addition to $\varepsilon$-factorization. The simplest iterated integrals found in Feynman integrals are multiple polylogarithms (MPLs) Goncharov:1998kja ; Goncharov:2001iea ; Vollinga:2004sn , which are iterated integrals over the genus-00 Riemann sphere with marked points, or say over moduli space ${\cal M}_{0,n}$ Bogner:2014mha . For MPL integral families, canonical differential equations can be achieved through rational and algebraic transformations alone, and several tools are available to assist in identifying such transformations Gituliar:2017vzm ; Prausa:2017ltv ; Meyer:2017joq ; Lee:2020zfb ; Dlapa:2020cwj . An alternative approach constructs canonical bases through ${\rm d}\log$-forms Henn:2020lye ; Chen:2020uyk ; Chen:2022lzr . [^2] However, not all integrals in ${\rm d}\log$-forms are polylogarithmic, see Duhr:2020gdd . MPLs exhibit properties that facilitate algorithmic extraction of critical information, aiding in the construction of canonical bases and enriching our understanding of their mathematical structure.
@@
Due to the success of the method of canonical differential equations in solving the MPL integral families, there have been substantial efforts to extend the method to the elliptic cases appearing in various scattering processes Muller:2022gec ; Jiang:2023jmk ; Gorges:2023zgv ; Delto:2023kqv ; Wang:2024ilc ; Forner:2024ojj ; Schwanemann:2024kbg ; Marzucca:2025eak ; Becchetti:2025rrz ; Becchetti:2025oyb . In such cases, there are subtleties in the precise definition of “canonical” Broedel:2018iwv ; Broedel:2018qkq ; Broedel:2019hyg ; Dlapa:2022wdu ; Frellesvig:2023iwr ; Duhr:2024uid ; Duhr:2025lbz , but one may at least attempt to obtain an $\varepsilon$-factorized system of differential equations which still offers simplification in the calculations. There are now several methods for deriving $\varepsilon$-factorized bases for elliptic cases vonManteuffel:2017hms ; Adams:2018bsn ; Frellesvig:2021hkr ; Pogel:2022vat ; Giroux:2022wav ; Dlapa:2022wdu ; Gorges:2023zgv , and some of them can even be extended to more complex geometries, such as Calabi-Yau varieties. In the elliptic cases, these methods usually construct pre-canonical bases according to the elliptic curves defined by some degree-3 or degree-4 polynomials, where other kinematic singularities appear as marked points. Further transformations are then applied to convert the differential equations for these pre-canonical bases to the $\varepsilon$-factorized form.
@@
It is well-known that the shapes of the elliptic curves depend only on the elliptic modulus or, equivalently, the modular lambda function $\lambda$. This is particularly evident in the so-called normal forms of elliptic curves, such as the Legendre normal form or the Jacobi normal form. In this paper, we employ this fact to systematize the construction of canonical bases for univariate elliptic Feynman integrals under the maximal cut. We use the term “canonical” here since it meets most of the important properties advocated by the literature Dlapa:2022wdu ; Frellesvig:2023iwr ; Duhr:2024uid . Our method utilizes the Legendre normal form of elliptic curves, where the geometric properties of the curves are simple and explicit. [^3] The Legendre normal form has also been employed in Broedel:2018rwm . Similarly, the Rosenhain normal form for the hyperelliptic case has been explored recently in Duhr:2024uid . We then construct the pre-canonical basis using Abelian differentials of three kinds and straightforwardly convert them to the canonical form with a universal transformation. This transformation is found by a method similar to Pogel:2022vat , and the result is also equivalent to Gorges:2023zgv . Due to the simplicity of the normal form, we are able to generically write down the canonical basis for a Legendre elliptic family with an arbitrary number of kinematic marked points. This generic canonical basis is the main new result of this paper. Finally, the Möbius transformation relating an arbitrary elliptic curve and its Legendre normal form can be written in a generic way as well. With such a transformation, our result can be directly applied to any univariate elliptic Feynman integral families.
@@
To demonstrate our construction, we apply the method to several concrete examples: fully massive sunrise and non-planar double box diagrams, each with two kinds of mass configurations. We show that it is straightforward to derive an $\varepsilon$-factorized system of differential equations within the top sectors. To obtain a canonical basis for the entire integral family, we still need to address the non-$\varepsilon$-factorized dependence on the sub-sectors. Ideally, one would imagine that it is possible to construct the full integrand without any cut [^4] We have demonstrated an example with the next-to-maximal cut in Sec. 3.2., combining the top-sector construction in this work and the ${\rm d}\log$-form construction in Chen:2020uyk ; Chen:2022lzr . In this work, we simply convert the differential equations to the $\varepsilon$-factorized form through a linear transformation. We show in the examples that the nice properties of our basis help a lot in this process.
@@
This paper is organized as follows. In Sec. 2, We present the construction of the canonical basis under maximal cut, which involves the construction in the Legendre normal form and the transformation to an elliptic curve defined by an arbitrary degree-4 polynomial. We demonstrate the procedure with several examples in Sec. 3, where the transformation to a canonical basis for the full family without cut is discussed. The summary and outlook are provided in Sec. 4. There are a few appendices at the end of the paper. A different strategy vonManteuffel:2017hms ; Frellesvig:2021hkr for deriving an $\varepsilon$-factorized basis is presented in App. A. The construction using the Jacobi normal form of elliptic curves is discussed in App. B. Some lengthy expressions relevant to the examples in Sec. 3 are collected in App. C. Special relations between elliptic integrals are compiled in App. D.
@@ 2.1 Basics for elliptic Feynman integrals
$I_{\nu_{1}\cdots\nu_{N}}=e^{L\varepsilon\gamma_{E}}(\mu^{2})^{\nu-Ld/2}\int\,\bigg{[}\,\prod_{j=1}^{L}\frac{{\rm d}^{d}k_{j}}{i\pi^{d/2}}\,\bigg{]}\frac{1}{D_{1}^{\nu_{1}}\cdots D_{N}^{\nu_{N}}}\,,$ (1)
@@
$u=u(z)=[P_{4}(z)]^{-1/2}\prod_{i=1}^{n}(z-c_{i})^{-\beta_{i}\varepsilon}\,,$ (3)
@@ 2.2 The construction of the pre-canonical basis
$u_{L}(x)=[P_{L}(x)]^{-1/2}\prod_{i=2}^{n+1}(x-e_{i})^{-\beta_{i}\varepsilon}\,,$ (13)
@@
${\rm d}\vec{I}=\bm{A}\cdot\vec{I}=\left({\bm{A}^{(0)}}+\varepsilon{\bm{A}^{(1)}}\right)\cdot\vec{I}\,,$ (19)
@@ 3.1.1 Definition
$u(z_{4})=[P_{4}(z_{4})]^{-1/2}\prod_{i=1}^{5}(z_{4}-c_{i})^{-\beta_{i}\varepsilon}\,.$ (62)
@@ 3.2.1 Definition
$s\equiv(p_{1}+p_{2})^{2},\quad t\equiv(p_{1}+p_{3})^{2}.$ (75)
@@
$u(z_{6},z)=[P_{4}(z)]^{-1/2}[P_{2}(z_{6},z)]^{-1/2}\prod_{i=1}^{4}(z-c_{i})^{-\beta_{i}\varepsilon}\prod_{j=6}^{7}(z_{6}-c_{j})^{-\beta_{j}\varepsilon}\,.$ (79)
@@ 3.2.3 Mixing with non-elliptic sectors
The $\varepsilon^{0}$-terms can be removed with regular “integrating out” strategy, and it turns out there are no new functions needed in the transformation beyond the periods $\varpi_{0}(\lambda)$ and $\vartheta_{0}(e_{5},\lambda)$ in ${\cal T}_{\rm can}$. For simplicity, we suppress the explicit expressions here, which can be found in ancillary files.
@@ 4 Conclusion and outlook
In this paper, we develop a systematic approach to constructing a canonical basis for elliptic Feynman integral families involving multiple kinematic scales, which exhibit a univariate integral in the Baikov representation under maximal cut. The main idea is to perform the construction using the Legendre normal form of elliptic curves, where the geometric information is explicit, and then apply a Möbius transformation to map it back to the original integral family. The pre-canonical basis integrals under maximal cut are expressed as linear combinations of Abelian differentials without elliptic integrals in the combination coefficients. This feature provides simplification when performing the $\varepsilon$ factorization for the dependence on sub-sectors. Furthermore, the canonical bases exhibit favorable behaviors around chosen degenerate points, where they asymptotically take the ${\rm d}\log$-forms. This helps to obtain simple boundary conditions for the differential equations. Around the degenerate point, only the dependence of the basis integrals on the marked points becomes significant, allowing the integrals to be expressed in terms of MPLs in this context.
@@
There can be different ways to construct the pre-canonical bases and the transformations bringing them to the canonical ones, and the approach we offer here is largely independent of the specific construction. Besides the method described in the main text, which is also equivalent to the one from Gorges:2023zgv , we have explored several other possibilities which are described in the appendices. In App. A, we show that our pre-canonical basis can be transformed to a canonical one by converting the period matrix into a constant matrix without $\varepsilon$-dependent transformation vonManteuffel:2017hms ; Frellesvig:2021hkr . Despite a simpler $\varepsilon$-dependence, this method introduces numerous elliptic integrals into the transformation, which, in turn, appear in the $\varepsilon$-factorized differential equations. In addition, this approach inherently leads to the transformation with logarithmic behaviors near degenerate points due to the complementary arguments $\lambda$ and $1-\lambda$ in elliptic integrals of the first and second kinds. Relevant discussions are included in Frellesvig:2023iwr . In contrast, the approach present in the main text avoids these issues, and the resulting canonical bases satisfy most of the nice properties advocated in Dlapa:2022wdu ; Frellesvig:2023iwr ; Duhr:2024uid .
@@
To demonstrate our method, we use fully massive sunrise families and non-planar double-box families with an internal mass as proof of concept. We emphasize again that, while our construction of the pre-canonical bases in the top-sectors and the transformation to canonical ones are equivalent to the methods in vonManteuffel:2017hms ; Adams:2018bsn ; Frellesvig:2021hkr ; Giroux:2022wav ; Dlapa:2022wdu ; Pogel:2022vat ; Gorges:2023zgv , these only need to be done once and for all, and the universal results have already been given in our paper. The applications to different integral families merely boil down to appropriate Möbius transformations. Moreover, since the proposed approach is not inherently limited to elliptic cases, it is expected to extend to more complex geometries, e.g., hyperelliptic cases, as explored in the recent work Duhr:2024uid . For example, using the canonical basis integrals constructed for the four-parameter Lauricella functions, one may derive the canonical basis for the equal-mass non-planar double (crossed) box family, at least under the maximal cut. One can also try constructing a canonical basis in the Rosenhain normal form family with kinematic marked points, which can be potentially applied for generic hyperelliptic Feynman integrals.
@@
Our method can be applied to integral families where multiple sectors involve the same elliptic curve. In such cases, the construction can be done sector-by-sector. However, we have found that it is helpful to perform the construction for these sectors together by loosening the maximal cut. This has been demonstrated for the non-planar double box family in Sec. 3.2, where both the top sector and the triangle sub-sector are elliptic. It remains possible that one may construct the full integrand without any cut, combining the elliptic-sector construction in this work and the ${\rm d}\log$-form construction in Chen:2020uyk ; Chen:2022lzr . It is also interesting to explore the application of our method to cases where multiple elliptic curves are involved. In this case, there can be non-trivial interplays among the Abelian differentials in different elliptic sectors, and it is interesting to see how these can be handled within our approach. Another complication arises in the boundary conditions since the degenerate points of the different elliptic curves may not coincide. It remains to be seen whether one may utilize the degenerate information separately, and connect them via the differential equations. We leave these for future investigations.